OpenStudy (anonymous):

12(x - 6) = 2(x - 9)

OpenStudy (saifoo.khan):

6x - 72 = x - 9

OpenStudy (anonymous):

12x-72=2x-18 10x=54 x=5.4

OpenStudy (saifoo.khan):

hey wait. 54/10

OpenStudy (saifoo.khan):

27/5 is the final answer.

OpenStudy (anonymous):

12x - 72 = 2x - 18 12x-2x = -18 + 72 10x = 54 x = 54/10 = 27/5